# Long-Horizon Model-Based Offline Reinforcement Learning Without Conservatism

### v1

@@ 4.4 Overall Algorithm
We now present the full method in Algorithm 1, integrating all design choices with the core rollout subroutine (Algorithm 2). Neubay remains conceptually simple, using single-step models and no uncertainty penalty, while carefully addressing the challenges of Bayesian offline RL. We highlight the lines where Neubay differs from prior work. For completeness, Appendix E provides details on RL loss and stopping criteria, although they are not needed to follow the experiments below.
@@ 5.2 What Matters in Neubay?
LayerNorm in the world model and uncertainty-based truncation jointly enables long-horizon rollouts. Figure 5 (see subsection H.2 for full results) shows that without LN, predicted state norms diverge quickly (around 10 steps here), driving exploding compounding errors. With LN, state norms remain bounded, which suppresses state error growth and stabilizes reward predictions. This matches our intuition based on Equation 5: by normalizing features at each step, LN constrains prediction magnitudes and thus compounding error. Moreover, the rightmost scatter plot shows what uncertainty-based truncation would accomplish: although we display full rollouts to reveal their divergence, applying thresholds would cut them off before entering high-error regions (here, around $\geq 10^{1}$). Thus, LayerNorm prevents error explosion, while uncertainty cutoff provides a complementary safeguard.
Adaptive long-horizon planning is the decisive factor for success. We study the role of adaptive horizons by varying $\zeta\in\{0.9,0.99,0.999,1.0\}$. Almost on all datasets, we find that the most aggressive choice, $\zeta=1.0$, delivers the best performance. Figure 1 shows a typical case, with full plots in subsection H.3, a summary in Table 2, and complete results in Table 13.
@@
On the other hand, performance often collapses to near zero (i.e., scores $\leq 5$) with smaller thresholds such as $0.9$ (16 datasets) or $0.99$ (6 datasets). The consistent failure mode behind these cases is severe value overestimation, where estimated Q-values ($\mathbb{E}_{(h_{t},a_{t})\sim\mathcal{D}}{\left[Q_{\omega}(h_{t},a_{t})\right]}$) are much higher (2nd column in Figure 1) compared to the actual performances. These results support our intuition in Equation 6: without conservatism, Bayesian RL mitigates overestimation by relying more on imagined rewards (lower-biased as shown in Figure 5 reward panel) using long horizons than on bootstrapped values. Crucially, adaptive horizons allow the algorithm to “trust” the model only within its in-distribution confidence region. As indicated by Figure 4, the region at $\zeta=1.0$ is much larger than at $\zeta=0.9$, yielding substantially longer horizons that provide useful data for both augmentation and mitigating overestimation.
@@
Introducing conservatism to Neubay helps some tasks, but not on average. To study conservatism in Neubay, we penalize the rewards with $\lambda\frac{U_{\boldsymbol{\theta}}(\hat{s},\hat{a})}{\mathbb{E}_{(s,a)\sim\mathcal{D}}{\left[U_{\boldsymbol{\theta}}(s,a)\right]}}$, normalized by the dataset average so that $\lambda$ is more comparable across datasets. As summarized in Table 2, a strong uncertainty penalty ($\lambda=1.0$ or $5.0$) generally hurts performance, while a small penalty ($\lambda=0.04$) performs comparably to Neubay ($\lambda=0$). The impact remains dataset-dependent, as detailed in Table 13.
@@
Context encoder learning rate and real data ratio have to be tuned per dataset. The best values of $\eta_{\phi}$ and $\kappa$ are reported in Table 4Table 6, with selective learning curves shown in Figure 13Figure 14. We find the optimal encoder learning rates are generally smaller than those used in online POMDPs (Luo et al., 2024a). For example, $1\text{\times}{10}^{-6}$ and $3\text{\times}{10}^{-7}$ yield the best results on 7 datasets. Our intuition is that, in Bayesian offline RL, smaller learning rates help curb overestimation by slowing down learning. For the real data ratio, we find $\kappa=0.05$, widely used in prior conservative algorithms (Yu et al., 2020), often yields poor performance. In Bayesian offline RL, large $\kappa$ acts as a softened uncertainty penalty, limiting overtrust in the model while avoiding explicit penalties.
Larger ensemble size improves performance. As summarized in Table 2 and detailed in Table 13, reducing the ensemble size $N$ to $20$ or $5$ clearly degrades performance, though not drastically on most tasks. This indicates that $N=100$ is close to the practical limit of what ensembling can offer for these tasks, and future work may explore more scalable ways to approximate the posterior.
Using a Markov agent in Neubay works well on most locomotion tasks, but fails in AntMaze. Popular offline RL methods employ Markov agents, while Neubay uses history-dependent agents due to its epistemic POMDP formulation. To assess the practical importance of memory, we replace the history-dependent agent in Neubay with a Markov one, sweep over the same real-data-ratio range, and report the best results in Table 2 and Table 13. The Markov version performs similarly to the history-dependent one on most locomotion tasks, especially on the NeoRL benchmark (64.7 $\to$ 66.8). However, it suffers severe degradation on hopper-medium-replay-v2 (110.6 $\to$ 47.2) and on all AntMaze tasks (43.2 $\to$ 1.3). This pattern suggests that in locomotion domains, epistemic uncertainty is relatively mild, so using a single observation may often infer the model index, making memory less critical. As a result, Neubay with a Markov agent remains a strong baseline in these settings. In contrast, AntMaze has high epistemic uncertainty, especially about the maze layout which is crucial to navigation, so memory is needed to accumulate information over time.
@@ 6 Conclusion and Future work
Tuning guidelines for Neubay. Drawing on our benchmarking and sensitivity studies in section 5, we suggest the following guidelines when applying Neubay to new tasks or after modifying a key module (e.g., world model, uncertainty quantifier, or memory encoder):
@@
  • Real data ratio. Adjust $\kappa$ with $[0.5,0.8]$ as a robust starting range.
  • Overestimation control. Monitor estimated values; if severe overestimation occurs, mitigate it by reducing the discount factor $\gamma$ or lowering $\eta_{\phi}$, as suggested by our analysis in Equation 6.
Conclusion. In this paper, we revisit conservatism as the dominant principle in offline RL and show that it is not universally optimal. Instead, we advance a Bayesian perspective that models epistemic uncertainty and trains history-dependent agents to maximize expected rewards over a posterior of world models. Building on this principle, we propose Neubay, a practical algorithm that mitigates compounding error with layer normalization, reduces value overestimation with adaptive long-horizon planning, and stabilizes recurrent training. Experiments across diverse benchmarks show that Neubay generally matches or surpasses strong conservative baselines, excels particularly on low-quality and moderate-coverage datasets, and challenges the belief that short-horizon planning is necessary in model-based RL.
@@ A.1 Offline Conservative Model-Based Methods
Our work differs from these prior lines of research in principle, core components, and algorithmic design. We replace conservatism with Bayesianism as the guiding principle, substituting the standard Markov actor-critic with a history-dependent one required by Bayesian principle, and derive the Neubay algorithm with adaptive long-horizon planning, fundamentally distinct from prior methods. Moreover, prior works typically rely on tuning two critical hyperparameters: the conservatism coefficient $\lambda$ and rollout horizon $H$, to achieve strong performance. In contrast, Neubay eliminates the need for $\lambda$ by design and replaces $H$ with an uncertainty quantile $\zeta$, which remains fixed at $1.0$ in main experiments.
@@ A.2 Offline Bayesian-Inspired and Non-Conservative RL
Similar to our algorithm, MAPLE (Chen et al., 2021c) and MoDAP (Choi et al., 2024) learn an ensemble of models for recurrent policy training. Unlike our approach, however, they store the hidden states of recurrent policies in the replay buffer and reuse them to initialize the policy’s context for rollouts and updates. This design is known to induce context staleness in recurrent RL (Kapturowski et al., 2018), whereas our algorithm avoids this issue by storing and sampling full histories directly. MAPLE also reintroduces an uncertainty penalty and terminates rollouts based on a predefined state boundary, making it a conservative method. MoDAP remains penalty-free but fine-tunes the world models during policy learning to maintain ensemble diversity; therefore, its objective departs from the epistemic POMDP, which requires freezing the posterior during policy optimization.
@@
In contrast, our work aims to stay as close as possible to the epistemic POMDP: we avoid conservatism and do not fine-tune models during policy learning. The approximation in the algorithm level arises in the rollout distribution: we begin rollouts from intermediate histories and truncate them using an uncertainty threshold to mitigate compounding errors. As discussed in subsection 4.2, using long rollouts keeps this deviation small in practice.
@@ A.3 Model-Generated Rollouts in RL
Mechanism of adaptive horizon. As discussed in subsection 4.2, several prior works have used uncertainty thresholds to adaptively truncate rollouts. In the online setting, Pan et al. (2020) truncates rollouts for states whose uncertainty ranks in the top 25% of the current-step batch and caps the horizon at $10$. Similarly, Frauenknecht et al. (2024) computes a 95% uncertainty quantile from the batch of first-step predictions in current rollouts, also with a maximum horizon of $10$. In the offline setting, Zhan et al. (2021) adopts an 85% uncertainty quantile from the offline dataset to filter rollouts and imposes a maximum horizon of $16$. Zhang et al. (2023) uses truncation thresholds proportional to the maximum uncertainty in the offline dataset, with a coefficient of $0.5$ and a maximum horizon of $5$. Overall, prior work employs adaptive horizons in a conservative manner, combining small thresholds with strict horizon caps. In contrast, our approach removes any maximum-horizon constraint and uses a large quantile threshold.
@@ Appendix B Formal Connections between Conservatism and Robustness
Appendix C Understanding on Bayesian Offline RL
@@ C.2 Diminishing Bootstrapped Error in Long-Horizon Rollouts
We formalize the insight in Equation 6 using a result adapted from Sims et al. (2024, Proposition 1).
@@
Thus, as long as the bootstrapped error $\epsilon$ dominates the intermediate TD errors, increasing rollout horizon $H$ exponentially suppresses the dominant source of error. In the model-based setting, these intermediate errors also include model prediction errors on rewards. Therefore, this formalizes our insight in Equation 6.
@@ Appendix G Implementation Details
G.1 World Model Ensemble and Planning
World model ensemble. Throughout our experiments, we train an ensemble of 128 MLPs on each dataset. Following MOPO (Yu et al., 2020), for continuous control tasks, we rank the models by MSE on a validation set consisting of 1000 held-out transitions from $\mathcal{D}$. For the bandit task, we rank by negative log-likelihood (NLL) that captures uncertainty, since the true reward follows Bernoulli distribution. The top $N$ models are then selected to form the world model ensemble $\mathbf{m}_{\boldsymbol{\theta}}$, which remains fixed during subsequent policy training. In our main experiments (subsection 5.1), we use $N=100$. For the sensitivity analysis (subsection 5.2), we vary this number to $5$ and $20$.
@@ G.1 World Model Ensemble and Planning
Following MOBILE (Sun et al., 2023; Sun, 2023) [^15] https://github.com/yihaosun1124/mobile., each MLP $m_{\theta}(s,a)$ has $4$ hidden layers with a width of $200$ for the locomotion benchmarks, and $5$ layers with a width of $400$ for Adroit benchmark. As described in subsection 4.1, we apply LayerNorm after each hidden layer; this design choice is further ablated in subsection 5.2. The basic block consists of (Linear $\to$ LayerNorm $\to$ leaky ReLU) when LayerNorm is enabled, and (Linear $\to$ leaky ReLU) when it is disabled. For the bandit task, we use a small reward-model ensemble that has $2$ hidden layers with a width of $16$. The weights of each MLP are initialized independently using the default equinox.nn.Linear scheme, i.e., $\text{Unif}[-\frac{1}{\sqrt{\text{dim}_{\text{in}}}},\frac{1}{\sqrt{\text{dim}_{\text{in}}}}]$, where $\text{dim}_{\text{in}}$ denotes the input feature dimension.
@@
For continuous control tasks, during ensemble training, we sample transition tuples $(s,a,r,s^{\prime})$ from $\mathcal{D}$ using a batch size of $256$ to estimate the MLE loss. The inputs $(s,a)$ and outputs $(r,s^{\prime})$ are standardized by subtracting the mean and dividing by the standard deviation computed over the dataset; during inference, predictions are inverse-transformed to restore the original scale. We use AdamW optimizer (Loshchilov & Hutter, 2019) with a weight decay coefficient of $5\text{\times}{10}^{-5}$, and a learning rate of $1\text{\times}{10}^{-3}$ for locomotion tasks and $3\text{\times}{10}^{-4}$ for Adroit tasks. Training is terminated early if the validation MSE fails to improve by more than $0.01$ relative within five consecutive epochs, following the early stopping procedure in MOBILE. In the bandit task, the model learning rate is $1\text{\times}{10}^{-3}$, batch size is $128$, and improvement threshold is $0.001$ absolute.

### v2

@@ 4.5 Overall Algorithm
We now present the full method in Algorithm 1, integrating all design choices with the core rollout subroutine (Algorithm 2). Neubay remains conceptually simple, using single-step models and no uncertainty penalty, while addressing the challenges of Bayesian offline RL. We highlight the lines where Neubay differs from prior work. In particular, to respect the Bayesian objective, we fix a single model per rollout, rather than randomizing the model index at each step as in MBPO-style methods. For completeness, Appendix E provides details on RL loss and stopping criteria, although they are not needed to follow the experiments below.
@@ 5.2 What Matters in Neubay?
Adaptive long-horizon planning (subsection 4.2) is the decisive factor for success. We study the role of adaptive horizons by varying $\zeta\in\{0.9,0.99,0.999,1.0\}$. Almost on all datasets, we find that the most aggressive choice, $\zeta=1.0$, delivers the best performance. Figure 1 shows a typical case, with full plots in subsection H.2, a summary in Table 2, and complete results in Table 13.
@@
On the other hand, performance often collapses to near zero (i.e., scores $\leq 5$) with smaller thresholds such as $0.9$ (16 datasets) or $0.99$ (6 datasets). The consistent failure mode behind these cases is severe value overestimation, where estimated Q-values ($\mathbb{E}_{(h_{t},a_{t})\sim\mathcal{D}}{\left[Q_{\omega}(h_{t},a_{t})\right]}$) are much higher (2nd column in Figure 1) compared to the actual performances. These results support our intuition in subsection 4.1: without conservatism, Bayesian RL mitigates overestimation by relying more on imagined rewards (lower-biased as shown in Figure 5 reward panel) using long horizons than on bootstrapped values. Crucially, adaptive horizons allow the algorithm to “trust” the model only within its in-distribution confidence region. As indicated by Figure 4, the region at $\zeta=1.0$ is much larger than at $\zeta=0.9$, yielding substantially longer horizons that provide useful data for both augmentation and mitigating overestimation.
@@
LayerNorm in the world model (subsection 4.3) and uncertainty truncation (subsection 4.2) jointly enables long-horizon rollouts. Figure 5 (see subsection H.3 for full results) shows that without LN, predicted state norms diverge quickly (around 10 steps here), driving exploding compounding errors. With LN, state norms remain bounded, which suppresses state error growth and stabilizes reward predictions. This matches our intuition based on Equation 6: by normalizing features at each step, LN constrains prediction magnitudes and thus compounding error. Moreover, the rightmost scatter plot shows what uncertainty-based truncation would accomplish: although we display full rollouts to reveal their divergence, applying thresholds would cut them off before entering high-error regions (here, around $\geq 10^{1}$). Thus, LayerNorm prevents error explosion, while uncertainty cutoff provides a complementary safeguard.
Larger ensemble size (subsection 4.3) improves performance. As summarized in Table 2 and detailed in Table 13, reducing the ensemble size $N$ to $20$ or $5$ clearly degrades performance, though not drastically on most tasks. This indicates that $N=100$ is close to the practical limit of what ensembling can offer for these tasks, and future work may explore more scalable ways to approximate the posterior.
Context encoder learning rate and real data ratio (subsection 4.4) have to be tuned per dataset. The best values of $\eta_{\phi}$ and $\kappa$ are reported in Table 4Table 6, with selective learning curves shown in Figure 13Figure 14. We find the optimal encoder learning rates are generally smaller than those used in online POMDPs (Luo et al., 2024a). For example, $1\text{\times}{10}^{-6}$ and $3\text{\times}{10}^{-7}$ yield the best results on 7 datasets. Our intuition is that, in Bayesian offline RL, smaller learning rates help curb overestimation by slowing down learning. For the real data ratio, we find $\kappa=0.05$, widely used in prior conservative algorithms (Yu et al., 2020), often yields poor performance. In Bayesian offline RL, large $\kappa$ acts as a softened uncertainty penalty, limiting overtrust in the model while avoiding explicit penalties.
Ablation: Introducing conservatism to Neubay helps some tasks, but not on average. To study conservatism in Neubay, we penalize the rewards with $\lambda\frac{U_{\boldsymbol{\theta}}(\hat{s},\hat{a})}{\mathbb{E}_{(s,a)\sim\mathcal{D}}{\left[U_{\boldsymbol{\theta}}(s,a)\right]}}$, normalized by the dataset average so that $\lambda$ is more comparable across datasets. As summarized in Table 2, a strong uncertainty penalty ($\lambda=1.0$ or $5.0$) generally hurts performance, while a small penalty ($\lambda=0.04$) performs comparably to Neubay ($\lambda=0$). The impact remains dataset-dependent, as detailed in Table 13.
@@
Ablation: Using a Markov agent in Neubay works well on most locomotion tasks, but fails in AntMaze. Popular offline RL methods employ Markov agents, while Neubay uses history-dependent agents due to its epistemic POMDP formulation. To assess the practical importance of memory, we replace the history-dependent agent in Neubay with a Markov one, sweep over the same real-data-ratio range, and report the best results in Table 2 and Table 13. The Markov version performs similarly to the history-dependent one on most locomotion tasks, especially on the NeoRL benchmark (64.7 $\to$ 66.8). However, it suffers severe degradation on hopper-medium-replay-v2 (110.6 $\to$ 47.2) and on all AntMaze tasks (43.2 $\to$ 1.3). This pattern suggests that in locomotion domains, epistemic uncertainty is relatively mild, so using a single observation may often infer the model index, making memory less critical. As a result, Neubay with a Markov agent remains a strong baseline in these settings. In contrast, AntMaze has high epistemic uncertainty, especially about the maze layout which is crucial to navigation, so memory is needed to accumulate information over time.
@@ 6 Conclusion and Future work
Tuning guidelines for Neubay. Based on our benchmarking and sensitivity studies, we suggest the following guidelines when applying Neubay to new tasks or after modifying key modules:
@@
  • Monitor overestimation. If severe overestimation occurs, mitigate it by reducing the discount factor $\gamma$ or lowering the learning rate $\eta_{\phi}$, as suggested by our analysis in subsection 4.1.
Conclusion. In this paper, we revisit conservatism as the dominant principle in offline RL and show that it is not universally optimal. Instead, we advance a Bayesian perspective that models epistemic uncertainty and trains history-dependent agents to maximize expected rewards over a posterior of world models. Building on this principle, we show that long-horizon rollouts play a critical role in reducing value overestimation once conservatism is removed, and we propose key design choices that enable both performing and learning from such rollouts, yielding the Neubay algorithm. Across diverse benchmarks, Neubay generally matches or surpasses strong conservative baselines, excels particularly on low-quality and moderate-coverage datasets, succeeds with long-horizon planning, and paves the way for a viable direction in offline and model-based RL.
@@ A.1 Offline Conservative Model-Based Methods
Our work differs from these prior lines of research in principle, core components, and algorithmic design. We replace conservatism with Bayesianism as the guiding principle, substituting the standard Markov actor-critic with a history-dependent one required by Bayesian principle, and derive the Neubay algorithm driven by long-horizon planning, fundamentally distinct from prior methods. Moreover, prior works typically rely on tuning two critical hyperparameters: the conservatism coefficient $\lambda$ and rollout horizon $H$, to achieve strong performance. In contrast, Neubay eliminates the need for $\lambda$ by design and replaces $H$ with an uncertainty quantile $\zeta$, which remains fixed at $1.0$ in main experiments.
@@ A.2 Offline Bayesian-Inspired and Non-Conservative RL
Similar to our algorithm, MAPLE (Chen et al., 2021c) and MoDAP (Choi et al., 2024) learn an ensemble of models for recurrent policy training. Unlike our approach, both algorithms perform short-horizon rollouts ($\leq 20$ steps). Moreover, both store the hidden states of recurrent policies in the replay buffer and reuse them to initialize the policy’s context for rollouts and updates. This design is known to induce context staleness in recurrent RL (Kapturowski et al., 2018), whereas our algorithm avoids this issue by storing and sampling full histories directly. MAPLE also reintroduces an uncertainty penalty and terminates rollouts based on a predefined state boundary, making it a conservative method. MoDAP remains penalty-free but fine-tunes the world models during policy learning to maintain ensemble diversity; therefore, its objective departs from the epistemic POMDP, which requires freezing the posterior during policy optimization. In contrast, our work aims to stay as close as possible to the epistemic POMDP: we avoid conservatism, develop design choices for long-horizon rollouts, and do not fine-tune models during policy learning.
@@
Test-time adaptation and offline-to-online RL. Finally, although our method performs test-time online adaptation, it does not involve online RL and therefore does not fall under the offline-to-online RL setting. The online adaptation arises purely from in-context memory, without any parameter updates. While one could view in-context memory as inducing an implicit policy improvement (Moeini et al., 2025), we do not treat this mechanism as explicit online learning in this work. In offline-to-online RL, Sentenac et al. (2025) analyzes bandit problems and shows that optimism can be preferable to pessimism during online learning, depending on the available interaction budget. While their focus is on a different regime, this observation is conceptually related to our finding that pessimism can hinder online adaptation.
@@ A.3 Model-Generated Rollouts in RL
Mechanism of adaptive horizon. As discussed in subsection 4.2, several prior works have used uncertainty thresholds to adaptively truncate rollouts. In the online RL setting, M2AC (Pan et al., 2020) truncates rollouts for states whose uncertainty ranks in the top 25% of the current-step batch and caps the horizon at $10$. Similarly, MACURA (Frauenknecht et al., 2024) computes a 95% uncertainty quantile from the batch of first-step predictions in current rollouts, also with a maximum horizon of $10$. Infoprop (Frauenknecht et al., 2025) extends MACURA by using accumulated uncertainty along the rollout as an additional truncation criterion and increases the maximum horizon to around $50$. In the offline RL setting, MOPP (Zhan et al., 2021) adopts an 85% uncertainty quantile from the offline dataset to filter rollouts and imposes a maximum horizon of $16$. TATU (Zhang et al., 2023) uses truncation thresholds proportional to the maximum uncertainty in the offline dataset, with a coefficient of $0.5$ and a maximum horizon of $5$.
Overall, prior work employs adaptive horizons in a conservative manner, combining small thresholds with strict horizon caps. In contrast, our approach removes any maximum-horizon constraint and uses a large quantile threshold.
@@ A.4 Bayesian and Partially Observable RL
A.5 Concurrent Work
ADM-v2 (Anonymous, 2026) is a concurrent work that, like ours, demonstrates the effectiveness of long-horizon rollouts for offline model-based RL. ADM-v2 pursues an extreme full-horizon setting, performing rollouts up to the maximum episode length (e.g., $1000$ steps in D4RL locomotion tasks) and thus treating the learned world model as a surrogate simulator. Building on ADMPO (Lin et al., 2025), it simplifies multi-step modeling and introduces a parallelized rollout mechanism, enabling efficient any-step prediction. With full-horizon rollouts, ADM-v2 can perform off-policy evaluation directly without value bootstrapping and achieves strong performance on popular benchmarks when combined with an uncertainty penalty, representing the first successful full-horizon method.
We view ADM-v2 as highly complementary to our work. While ADM-v2 demonstrates that full-horizon rollouts can be effective under a conservative design, our method explores the opposite extreme: adhering to a Bayesian, non-conservative objective while using adaptive (but not full) horizons. Interestingly, ADM-v2’s sensitivity analysis (Anonymous, 2026, Figure 16) shows that long horizons are most beneficial under strong penalties, whereas in lightly penalized regimes, excessively long horizons can degrade performance. This aligns with our findings: without conservatism, performance is more sensitive to compounding errors, motivating adaptive truncation rather than full-horizon rollouts. Together, these two works suggest that long-horizon planning is a powerful component in offline RL, whose effective use depends on the degree of conservatism.
@@ Appendix B Formal Connections between Conservatism and Robustness
Appendix C Understanding of Bayesian Offline RL
@@ C.2 Diminishing Bootstrapped Error in Long-Horizon Rollouts
We formalize the insight in subsection 4.1 using a result adapted from Sims et al. (2024, Proposition 1).
@@
The bound above is sign-agnostic, but under offline policy improvement the bootstrapped term at the truncation point tends to be optimistic. Near-greedy policies evaluate the critic on out-of-distribution actions that require extrapolation, where neural critics can assign spuriously high values. Even though edge-of-reach histories are not directly updated, their values can drift through function approximation and shared parameters, allowing optimistic errors to propagate. As a result, the terminal bootstrap often dominates value overestimation for short horizons, while increasing $H$ mitigates this effect by exponentially down-weighting it.
@@ Appendix G Implementation Details
G.1 Details on World Model Ensemble
We provide implementation details on world ensemble, as introduced in section 2 and subsection 4.3.
@@ G.1 Details on World Model Ensemble
Ensemble architecture. Following common practice in offline RL (Yu et al., 2020), our world ensemble is a set of neural networks $\mathbf{m}_{\boldsymbol{\theta}}=\{m_{\theta^{n}}\}_{n=1}^{N}$ where each model outputs a Gaussian distribution over next state $s^{\prime}$ and reward $r$: $m_{\theta^{n}}(s,a)=\mathcal{N}(\mu_{\theta^{n}}(s,a),\sigma_{\theta^{n}}(s,a))$, with parameters $\boldsymbol{\theta}=\{\theta^{n}\}_{n=1}^{N}$ independently initialized at random. Following MOBILE (Sun et al., 2023; Sun, 2023) [^15] https://github.com/yihaosun1124/mobile., each MLP $m_{\theta}(s,a)$ has $4$ hidden layers with a width of $200$ for the locomotion benchmarks, and $5$ layers with a width of $400$ for Adroit benchmark. As described in subsection 4.3, we apply LayerNorm after each hidden layer; this design choice is further ablated in subsection 5.2. The basic block consists of (Linear $\to$ LayerNorm $\to$ leaky ReLU) when LayerNorm is enabled, and (Linear $\to$ leaky ReLU) when it is disabled. For the bandit task, we use a small reward-model ensemble that has $2$ hidden layers with a width of $16$. The weights of each MLP are initialized independently using the default equinox.nn.Linear scheme, i.e., $\text{Unif}[-\frac{1}{\sqrt{\text{dim}_{\text{in}}}},\frac{1}{\sqrt{\text{dim}_{\text{in}}}}]$, where $\text{dim}_{\text{in}}$ denotes the input feature dimension.
Ensemble selection. For each dataset, we train an initial pool of 128 MLPs and select the top $N$ models to form the world model ensemble $\mathbf{m}_{\boldsymbol{\theta}}$, which remains fixed during subsequent policy training. Model selection is based on a held-out validation set. For continuous control tasks, we rank the models by MSE on a validation set consisting of 1000 transitions from $\mathcal{D}$, following MOPO (Yu et al., 2020). For the bandit task, we rank by negative log-likelihood (NLL) that captures uncertainty, since the true reward follows Bernoulli distribution. In our main experiments (subsection 5.1), we use $N=100$. For the sensitivity analysis (subsection 5.2), we vary $N\in\{5,20\}$.
@@
Ensemble training. Each model is trained using standard maximum likelihood estimation (MLE). For continuous control tasks, we sample transition tuples $(s,a,r,s^{\prime})$ from $\mathcal{D}$ using a batch size of $256$ to estimate the MLE loss. The inputs $(s,a)$ and outputs $(r,s^{\prime})$ are standardized by subtracting the mean and dividing by the standard deviation computed over the dataset; during inference, predictions are inverse-transformed to restore the original scale. We use AdamW optimizer (Loshchilov & Hutter, 2019) with a weight decay coefficient of $5\text{\times}{10}^{-5}$, and a learning rate of $1\text{\times}{10}^{-3}$ for locomotion tasks and $3\text{\times}{10}^{-4}$ for Adroit tasks. Training is terminated early if the validation MSE fails to improve by more than $0.01$ relative within five consecutive epochs, following the early stopping procedure in MOBILE. In the bandit task, the model learning rate is $1\text{\times}{10}^{-3}$, batch size is $128$, and improvement threshold is $0.001$ absolute.
G.2 Details on Planning